### v1

@@ Abstract
The recent advance in video understanding has been driven by multimodal large language models (MLLMs). But these MLLMs are good at analyzing short videos, while suffering from difficulties in understanding videos with a longer context. To address this difficulty, several agent paradigms have recently been proposed, using MLLMs as agents for retrieving extra contextual knowledge in a long video. However, most existing agents ignore the key fact that a long video is composed with multiple shots, i.e., to answer the user question from a long video, it is critical to deeply understand its relevant shots like human. Without such insight, these agents often mistakenly find redundant even noisy temporal context, restricting their capacity for long video understanding. To fill this gap, we propose VideoChat-A1, a novel long video agent paradigm. Different from the previous works, our VideoChat-A1 can deeply think with long videos, via a distinct chain-of-shot reasoning paradigm. More specifically, it can progressively select the relevant shots of user question, and look into these shots in a coarse-to-fine partition. By multi-modal reasoning along the shot chain, VideoChat-A1 can effectively mimic step-by-step human thinking process, allowing to interactively discover preferable temporal context for thoughtful understanding in long videos. Extensive experiments show that, our VideoChat-A1 achieves the state-of-the-art performance on the mainstream long video QA benchmarks, e.g., it achieves 77.0 on VideoMME and 70.1 on EgoSchema, outperforming its strong baselines (e.g., Intern2.5VL-8B
and InternVideo2.5-8B), by up to 10.8% and 6.2%. Compared to leading close-source GPT-4o and Gemini 1.5 Pro, VideoChat-A1 offers competitive accuracy, but with 7% input frames and 12% inference time on average. We will release the codes afterwards.
1 Introduction
@@
Video understanding is an important problem in computer vision [54], which has great potential in real-world applications such as surveillance, smart phones, robotics, etc. With fast development of multimodal large language models (MLLMs), video understanding has achieved significant progress [34, 41, 44, 6, 1, 62]. By aligning video encoder with LLM, these models can flexibly interact with videos in dialogues. However, most existing MLLMs are good at understanding short videos in seconds, while hardly analyze longer videos in minutes (or hours) properly. The main challenge lies in feeding a huge amount of frames in long videos to MLLMs. To deal with this problem, several attempts have been proposed by modeling such long multimodal context [34, 41, 27], or equipping MLLMs with token compression [24, 25, 39], But their efficiency and effectiveness still need to be further improved in tackling redundant video content, thus blocking their performance on long video understanding benchmarks.
@@
Figure 1: Motivation. Direct reasoning models such as GPT-4o [34] and Qwen2.5VL-72B [1] perform global sampling on videos and struggle to focus on key information in long videos. Agent-based methods like VideoTree [46] often suffer from incorrect or redundant sampling to generate noisy caption for wrong answer, due to the lack of deep thinking on shots in a long video. In contrast, VideoChat-A1 interactively employs shot perception and reasoning via Chain-of-Shot, which progressively looks into relevant shots through a reflective process to achieve superior performance.
@@
Recent studies have shown that, agent-based approaches are promising for long video understanding  [43, 4, 9, 18, 28, 52, 30, 2, 60, 46]. Instead of feeding the entire long video with model adaptation, they simplify this challenge by invoking various tools to extract and retrieve relevant information from long videos. However, these approaches lack flexibility to capture complex content changes in a long video, since the retrieval is based on off-the-shelf video knowledge which is fixed after extraction. Alternatively, OpenAI o3 [35] shows a human-like thinking process on images. Via chain-of-thought reasoning, o3 can deeply understand visual content via progressive interaction with images. A similar insight has been also introduced in understanding long videos [43, 46], i.e., instead of watching a long video once, these approaches iteratively find relevant video frames for question answering. However, they ignore that, a long video consists of multiple shots. To answer the user question reliably, it is critical to deeply understand its relevant shots via multi-round reasoning. Without this consideration, these agents often mistakenly find redundant even noisy temporal context, restricting their capacity for long video understanding.
@@
To fill this gap, we propose VideoChat-A1, a novel video agent framework that can progressively think with a long video, via an interactive chain-of-shot reasoning paradigm. Specifically, for each shot reasoning step, we leverage MLLM as core agent of thinking, and invoke tools for selecting relevant shots, dividing these shots into subshots, and reasoning user answer with these subshots. If the answer is unconfident, it means that the current subshots are insufficient for understanding user question. As a result, VideoChat-A1 will start next-step shot reasoning, in order to further understand user question with finer relevant shots. Through such a distinct chain-of-shot paradigm, our VideoChat-A1 can effectively mimic the thinking process of human, by progressively reasoning on user answer while iteratively looking into relevant video shots, as shown in Fig. 1. To show its effectiveness, we implement VideoChat-A1 on 3 strong MLLMs with 4 mainstream long video question answering benchmarks. Extensive experiments show that, our VideoChat-A1 achieves the state-of-the-art performance, e.g., it achieves 77.0 on VideoMME with subset and 70.1 on EgoSchema, outperforming its baselines such as InternVL2.5-8B and InternVideo2.5-8B, by up to 10.8% and 6.2%. Compared to GPT-4o and Gemini 1.5 Pro, VideoChat-A1 offers competitive accuracy, but with 7% input frames and 12% inference time on average.
2 Related Work
@@
Multimodal Large Models (MLLMs) have demonstrated great potential in the field of video understanding [34, 41, 58, 42, 22]. In tasks such as video question-answering, captioning, and grounding, MLLMs have significantly enhanced the understanding performance by constructing cross-modal semantic representations [25, 12, 40, 1, 6]. Despite the encouraging progress of MLLMs, they still face challenges when dealing with long-form videos that can be minutes or even hours in length and contain a substantial amount of redundant information [40]. Currently, there are two main approaches to addressing the problem of long-form videos. One is to compress visual tokens [40, 57, 47, 25, 10, 55]. For example, MA-LLM [12] merges similar tokens, and LLaMA-VID [25] compresses each frame into context and content tokens to enable longer video input. However, these methods result in a significant loss of visual information. Another approach is to extend the number of processable tokens [26, 41, 42, 51]. Models like LongVILA [51, 14, 24] have adopted strategies to increase the token capacity of LLMs (Large Language Models) for handling longer videos. However, this method introduces substantial redundant information, consumes more computational resources, and often achieves only moderate effectiveness.
@@
Agent mechanisms [21, 8, 45, 16, 38, 49, 3, 4, 18, 28, 53] have been gradually introduced in video understanding tasks to enhance the semantic modeling and information acquisition capabilities for MLLMs. Currently, Agent methods for long video understanding can be mainly divided into two categories. One is extracting knowledge from long videos by invoking external tools at once and answering questions by retrieving the extracted knowledge. For example, CLIP [36] is used for retrieving key frames related to the questions [59, 50]. Memory banks [9, 12] and search engines [20, 5] are applied for extracting key information of the videos. However, this single-pass information extraction approach may lead to the omission of critical details or the inclusion of excessive redundant information, both of which interfere with the model’s response. Some methods convert the long video into long text for understanding [5, 46, 43]. The other approach involves answering questions through a round by round key information search. The multi-round video understanding pipeline emphasizes active video exploration and dynamic information acquisition by the model [46, 43, 30], and optimizes the reasoning path through a multi-round retrieval and reflection mechanism. However, a critical limitation is the lack of focus on video shots, the fundamental structural units of long-form videos, during content search. This oversight can introduce noisy information or abrupt content shifts, thereby degrading model performance.
@@
3 Method
In this section, we introduce our VideoChat-A1 in detail. To deeply understand the user question in a long video, we design a distinct Chain-Of-Shot (CoS) reasoning paradigm, which can progressively think while discovering relevant shots via multi-round shot partition and dialogue with MLLMs.
Specifically, it consists of three key steps in each round of shot reasoning, including Shot Selection, Shot Partition, and Shot Reflection. The framework of VideoChat-A1 is shown in Fig. 2. We will explain all of them in the following.
@@ 3 Method
3.1 Shot Selection
For conciseness, we describe the $i$-th shot reasoning step for illustration. Suppose that, we divide the whole video into $M$ shots at the $(i-1)$-th round,
@@ 3.1 Shot Selection
where we treat the entire video as one single shot at the 1st round. As shown in Fig. 2, we first need to select the candidate shots from the existing ones at the $(i-1)$-th round, in order to further look into these candidates for reasoning at the $i$-th round.
@@
where $\mathcal{Q}$ and $\mathcal{O}$ refers to the user question and answer options. $\mathcal{V}^{(i-1)}$ refers to the video frames sampled from the shots in the historical rounds, e.g., we sample 32 frames from each selected shot at round 1. In each subsequent round, we uniformly sample 8 frames from each selected shot and add them into the frame set.
Moreover, $\mathcal{H}^{(i-1)}$ refers to historical reasoning information, including key text information $\mathcal{I}^{(i-1)}$, the chosen answer $\mathcal{A}^{(i-1)}$, and the reason why to choose this answer $\mathcal{R}^{(i-1)}$ at round $(i-1)$.
@@
where we leverage our finetuned LongCLIP [56] as a retriever,
and compute the cosine similarities between key information and each shot in $\mathcal{S}^{(i-1)}$. Finally, we select top $N$ shots as candidates for further investigation, $\mathcal{C}^{(i)}=\{\mathcal{C}_{1}^{(i)},...\mathcal{C}_{N}^{(i)}\}$, where $\mathcal{C}_{n}^{(i)}$ refers to the $n$-th selected shot from the shot set $\mathcal{S}^{(i-1)}$. Additionally, since there is one single shot (i.e., the whole video) at the 1st round, we directly use this shot as the candidate at this round.
3.2 Shot Partition
Figure 3: Shot Partition. Given a candidate shot at step $i$, VideoChat-A1 first applies K-Means clustering to obtain K cluster centers for finding key frames. Subsequently, subshots are partitioned based on the feature distance between each frame and its adjacent two key frames.
After obtaining the candidate shots $\mathcal{C}^{(i)}$, we next further divide each of them into subshots, in order to look deeper into it for reasoning. In this paper, we introduce a concise shot partition pipeline as shown in Fig. 3. For notation simplicity, we illustrate how to perform it on a candidate shot.
Key Frame Discovery. To discover the subshots in a candidate shot, we start by finding the key frames in the subshots. First, we uniformly sample frames from the candidate shot. To reduce temporal redundancy while maintaining computation efficiency, we sample 1 frame per second in this paper. Second, we extract the features of the sampled frames by CLIP. Third, we perform K-Means Clustering  [31] on these features. As a result, each of $K$ clusters roughly reflects a subshot. However, frames in each cluster may not be temporally adjacent, since K-Means Clustering does not take temporal order into account. To contruct subshots in a temporal order, we further find a key frame in each cluster to divide the shot. Specifically, for each cluster, we choose the frame whose feature is closest to the cluster center, as the key frame in this cluster. Finally, we organize $K$ key frames in the temporal order, $\mathcal{K}=\{\mathcal{K}(1),...,\mathcal{K}(K)\}$, for the candidate shot.
@@ 3.2 Shot Partition
3.3 Shot Reflection
Figure 4: Shot Reasoning and Shot Reflection. At each step, VideoChat-A1 performs question-answering reasoning using the relevant shots identified. It then evaluates the generated answer and historical information to reflect on the confidence level of the response. Based on the confidence score and the number of reasoning iterations, the system either proceeds to the next step for further refinement or terminates the reasoning process to output the final answer. This iterative reflection ensures reliable and contextually accurate responses.
Question Reasoning. After obtaining subshots from the candidate shots, we employ MLLMs to answer the user question based on these finer shot regions. Specifically, as shown in Fig. 4, we leverage MLLM as dialogue agent for answering question at round $i$,
@@ 3.3 Shot Reflection
where $\mathcal{Q}$ and $\mathcal{O}$ refers to question and answer options. Moreover, we sample 8 frames from each subshot $\mathcal{S}_{n}^{(i)}(k)$, and combine them with $\mathcal{V}^{(i-1)}$ to form $\mathcal{V}^{(i)}$. If some sampled frames are the same as those in $\mathcal{V}^{(i-1)}$, we would merge them and resample more frames to ensure frame diversity in $\mathcal{V}^{(i)}$. Additionally, note that, we treat the whole video as the candidate shot at the 1st round without selection. If we sample frames from each subshot from this candidate, we actually look globally for question answering. Hence, we leverage key text information at the 1st round as guidance, and use CLIP to retrieve the most relevant subshot for question answering at the 1st round. Finally, given these inputs, we ask MLLM to generate answer $\mathcal{A}^{(i)}$, and the reason why to choose this answer $\mathcal{R}^{(i)}$.
@@
In this work, we set the confidence level $\mathcal{Z}^{(i)}$ from 0 to 3. If the answer confidence is higher than 2, we believe that it is a confident answer as the final output. Otherwise, we believe that it is a unreliable answer, i.e., the found subshots at this round are not sufficient to answer the question. We need to further look into them to get the answer. Hence, we start shot selection, partition, reflection in the next step. Additionally, we set a maximum number of shot reasoning steps. If VideoChat-A1 still do not get the confident answer at the max step, it will vote the major answer as the final one, according to the answers of all the rounds. Via exploit relevant shots in a coarse-to-fine manner, our VideoChat-A1 allows to leverage a chain of shots for deep thinking on long videos, and thus progressively and interactively discover the reliable shot context to answer user question.
4 Experiment
@@
Close-Source Model
GPT-4o(0513) [34] 72.2 66.7 64.6 80.0 / 82.8 70.3 / 76.6 65.3 / 72.1 71.9 / 77.2
Gemini 1.5Pro [41] 71.1 64.0 - 81.7 / 84.5 74.3 / 81.0 67.4 / 77.4 75.0 / 81.3
@@
LLaVA-OneVision-72B [19] 62.0 63.2 68.0 76.7/79.3 62.2/66.9 60.0/62.4 66.3/69.6
VideoLLaMA-2-72B [7] 63.9 - 45.6 69.8/72.0 59.9/63.0 57.6/59.0 62.4/64.7
LLAVA-Video-72B [58] 65.6 64.9 - 81.4/82.8 68.9/75.6 61.5/72.5 70.6/76.9
Qwen2.5-VL-72B [1]  77.9 - - 80.1/82.2 71.3/76.8 62.2/74.3 71.2/77.8
InternVL-2.5-78B [6] - 63.6 75.7 82.8/83.2 70.9/74.1 62.6/64.8 72.1/74.0
InternVL-3-78B [62] - 65.7 79.5 - - - 72.7 /75.7
@@
ShareGPT4Video-8B [5] - 39.7 46.4 48.3/53.6 36.3/39.3 35.0/37.9 39.9/43.6
VideoChat2-7B [23] 56.7 39.3 47.9 48.3/52.8 37.0/39.4 33.2/39.2 39.5/43.8
LongVU-7B [37] 67.6 - 65.4 - - -/59.5 -/60.6
LLaVA-Video-7B [58] 57.3 58.2 70.8 - - - 63.3/69.7
Qwen2.5-VL-7B [1] 66.7 55.6 - - - - 63.3/69.0
InternVideo-2.5-8B [44] 63.9 60.6 72.8 - - - 65.1/-
InternVL-2.5-8B [6] - 60.0 68.9 - - - 64.2/66.9
InternVL-3-8B [62] - 58.8 71.4 - - - 66.3/68.9
@@
VideoAgent [43] 54.1 - - - - - -
VideoTree (GPT-4)  [46] 61.1 - - - - - -
DrVideo [30] 61.0 - - - - - 51.7/71.7
T* [53] - - - 61.0/- 66.6/- 77.5/- 68.3/-
VideoMind [28] - - 64.4 - - 49.2/- 58.2/-
VideoMultiAgents [18] 68.0 - - - - - -
VideoRAG-72B [29] - 65.4 73.8 81.1/- 72.9/- 73.1/- 75.7/-
@@
Table 1: Comparison with close-source, open-source, and agent based model on four Long Video Understanding Tasks. The Video-MME results are presented in the format “w/o subs / w/ subs".
@@
GPT-4o[34] 384 134.4s 72.2 66.7 64.6 71.9/77.2
Gemini-1.5-Pro[41] 568 198.8s 71.1 64.0 - 75.0/81.3
Qwen2.5-VL-72B[1] 768 122.4s 77.9 - - 71.2/77.8
Qwen2.5-VL-7B[1] 512 24.3s 66.7 55.6 - 63.3/69.7
InternVL-2.5-8B[6] 64 12.4s - 60.0 68.9 64.2/66.9
InternVideo-2.5-8B[44] 512 33.8s 63.9 60.6 72.8 65.1/-
@@
Table 2: Comparsion on Average Frame Number and Inference Time on Datasets. Inference Time includes the average time for video frame extraction, clustering, and the time required for the MLLM to answer all queries. Compared to closed-source or 72B/78B models, our method achieves competitive performance with shorter inference time and fewer frames. Compared to 7B/8B models, our method achieves significant performance improvements with less or comparable inference time.
4.1 Experimental Setup
Tasks & Datasets.
We test our VideoChat-A1 on four diverse long-form video question-answering benchmarks. EgoSchema [32] collects 5,032 multiple-choice question-answer pairs that span 250 hours of video content. The average video duration is 180s from an ego perspective. LongVideoBench [48] consists of 1,337 multiple-choice questions across 17 fine-grained categories average duration is 473 seconds. MLVU [61] includes 2,174 QA-pairs with an average video length of 15 minutes, ranging from 3 minutes to over 2 hours. Video-MME [11] includes 2,700 QA pairs of 900 videos. The videos’ length ranges from 3 minutes to 1 hour.
Implementation Details.
We adopt Qwen2.5-VL-7B, InternVL2.5-8B, and InternVideo2.5-8B as our baseline MLLMs.

### v2

@@ Abstract
Recent advances in video understanding have been driven by MLLMs. But these MLLMs are good at analyzing short videos, while suffering from difficulties in understanding videos with a longer context. To address this difficulty, several agent methods have been proposed, using MLLMs as agents for retrieving extra contextual knowledge in a long video. However, most existing agents ignore the key fact that a long video is composed with multiple shots, i.e., to answer the user question from a long video, it is critical to deeply understand its relevant shots like human. Without such insight, these agents often mistakenly find redundant even noisy temporal context, restricting their capacity for long video understanding. To fill this gap, we propose VideoChat-A1, a novel long video agent paradigm. Different from the previous works, our VideoChat-A1 can deeply think with long videos, via a distinct chain-of-shot reasoning paradigm. More specifically, it can progressively select the relevant shots of user question, and look into these shots in a coarse-to-fine partition. By multi-modal reasoning along the shot chain, VideoChat-A1 can effectively mimic step-by-step human thinking process, allowing the interactive discovery of preferable temporal context for thoughtful understanding in long videos. Extensive experiments show that, VideoChat-A1 achieves the state-of-the-art performance on the mainstream long video QA benchmarks, e.g., it achieves 77.0 on VideoMME (w/ subs) and 70.1 on EgoSchema, outperforming its strong baselines (e.g., InternVL2.5-8B and InternVideo2.5-8B), by up to 10.1% and 6.2%. Compared to leading closed-source GPT-4o and Gemini 1.5 Pro, VideoChat-A1 offers competitive accuracy, but only with 7% input frames and 12% inference time on average. The code is available on https://github.com/SpXace/VideoChat-A1.
Introduction
@@
Video understanding is an important problem in computer vision (You et al. 2024). With fast development of MLLMs, video understanding has achieved significant progress (OpenAI 2024; Wang et al. 2025). However, most existing MLLMs are good at understanding short videos in seconds, while hardly analyzing longer videos in minutes (or hours) properly. The main challenge lies in feeding a huge amount of frames in long videos to MLLMs. To deal with this problem, several attempts have been proposed by modeling such long multimodal context (OpenAI 2024; Liu et al. 2025a), or equipping MLLMs with token compression (Li et al. 2025a; Shu et al. 2024), But their efficiency and effectiveness still need to be further improved in tackling redundant video content, thus blocking their performance on long video understanding benchmarks.
@@
Figure 1: Motivation. Direct reasoning models such as GPT-4o (OpenAI 2024) perform global sampling and struggle to focus on key information. Agent-based methods like VideoTree (Wang et al. 2024c) often suffer from incorrect or redundant sampling that leads to noisy captions. In contrast, VideoChat-A1 interactively employs shot perception and reasoning via Chain-of-Shot, which progressively looks into relevant shots through a reflective process for better performance.
@@
Recent studies have shown that, agent-based approaches are promising for long video understanding  (Wang et al. 2024a; Chen et al. 2025; Kugo et al. 2025; Zhi et al. 2025; Wang et al. 2024c). Instead of feeding the entire long video with model adaptation, they simplify this challenge by invoking various tools to extract and retrieve relevant information from long videos. However, these approaches lack flexibility to capture complex content changes in a long video, since the retrieval is based on off-the-shelf video knowledge which is fixed after extraction. Alternatively, OpenAI o3 (OpenAI et al. 2025) shows a human-like thinking process on images. Via chain-of-thought reasoning, o3 can deeply understand visual content via progressive interaction with images. However, they ignore that, a long video consists of multiple shots. To answer the user question reliably, it is critical to deeply understand its relevant shots via multi-round reasoning. Without this consideration, these agents often mistakenly find redundant even noisy temporal context, restricting their capacity for long video understanding.
@@
To fill this gap, we propose VideoChat-A1, a novel video agent framework that can progressively think with a long video, via an interactive chain-of-shot reasoning paradigm. Specifically, for each shot reasoning step, we leverage MLLM as core agent of thinking, and invoke tools for selecting relevant shots, dividing these shots into subshots, and reasoning answer with subshots. If the answer is unconfident, it means that the current subshots are insufficient for understanding user question. As a result, VideoChat-A1 will start next-step shot reasoning to further understand user question with finer relevant shots. Through such a distinct chain-of-shot paradigm, our VideoChat-A1 can effectively mimic the thinking process of human, by progressively reasoning on user answer while iteratively looking into relevant video shots, as shown in Fig. 1. Extensive experiments show that, our VideoChat-A1 achieves the state-of-the-art performance on 4 mainstream benchmarks, e.g., it achieves 77.0 on VideoMME with subset and 70.1 on EgoSchema, outperforming its baselines such as InternVL2.5-8B and InternVideo2.5-8B, by up to 10.1% and 6.2%. Compared to GPT-4o and Gemini 1.5 Pro, VideoChat-A1 offers competitive accuracy, but with 7% input frames and 12% inference time on average.
Related Work
@@
MLLMs have demonstrated great potential in the field of video understanding (OpenAI 2024; Li et al. 2023; Chen et al. 2022). However, they still face challenges when dealing with long videos that can be minutes or hours in length  (Song et al. 2024). Currently, there are two main approaches to addressing this problem. One is to compress visual tokens (Song et al. 2024; Zeng et al. 2024). For example, MA-LLM (He et al. 2024) merges similar tokens, and LLaMA-VID (Li et al. 2025b) compresses each frame into context and content tokens to enable longer video input. However, these methods result in a significant loss of visual information. Another approach is to extend the number of processable tokens (Liu et al. 2024; Wang et al. 2024b). Models like LongVILA (Xue et al. 2024; Li et al. 2025a) have adopted strategies to increase the token capacity for handling longer videos. However, this method introduces redundant information, consumes more computational resources, and often achieves only moderate effectiveness.
@@
Agent mechanisms (Chen et al. 2025; Kugo et al. 2025; Ye et al. 2025) have been introduced in video understanding tasks. Currently, Agent methods for long video understanding can be mainly divided into two categories. One is extracting knowledge from long videos by invoking external tools at once and answering questions by retrieving the extracted knowledge. For example, CLIP (Radford et al. 2021) is used for retrieving key frames related to the questions (Zhao et al. 2024; Xie et al. 2023). Memory banks (Fan et al. 2024; He et al. 2024) and search engines (Li et al. 2024b; Chen et al. 2024a) are applied for extracting key information of the videos. However, this single-pass information extraction often misses critical details or introduces redundant information, which interferes with the model’s response. The other approach involves answering questions through a round by round key information search. The multi-round pipeline emphasizes active video exploration and dynamic information acquisition (Wang et al. 2024c, a; Ma et al. 2024), and optimizes the reasoning path through retrieval and reflection mechanism. However, a critical limitation is that most methods either ignore video shots—the fundamental structural units of long-form videos—or segment shots at once without further fine-grained partitioning, which can both miss critical query-relevant details within each shot and introduce redundant or noisy information. For instance, VideoINSTA (Liao et al. 2024) proposes an event-based spatial-temporal reasoning framework, but its spatial and temporal information extraction processes are relatively independent and lack query-driven interactive iterative refinement.
@@
Method
In this section, we introduce our VideoChat-A1. To deeply understand the user question in a long video, we design a distinct Chain-of-Shot (CoS) reasoning paradigm, which can progressively think while discovering relevant shots via multi-round shot partition and dialogues. Specifically, it consists of three key steps in each round of shot reasoning, including Shot Selection, Shot Partition, and Shot Reflection. The framework of VideoChat-A1 is shown in Fig. 2.
@@
Shot Selection
For conciseness, we describe the $i$-th shot reasoning step for illustration. Suppose that, we divide the whole video into $M$ shots at the (i1)(i-1)-th round,
@@
As shown in Fig. 2, we first select the candidate shots from the existing ones at the (i1)(i-1)-th round, in order to further look into these candidates for reasoning at the $i$-th round.
@@
where $\mathcal{Q}$ and $\mathcal{O}$ refers to the user question and answer options. $\mathcal{V}^{(i-1)}$ refers to the video frames sampled from the shots in the historical rounds. Moreover, $\mathcal{H}^{(i-1)}$ refers to historical reasoning information, including key text information $\mathcal{I}^{(i-1)}$, the chosen answer $\mathcal{A}^{(i-1)}$, and the reason why to choose this answer $\mathcal{R}^{(i-1)}$ at round (i1)(i-1).
@@
where we leverage our finetuned LongCLIP (Zhang et al. 2024a) as a retriever, and compute the cosine similarities between key information and each shot in $\mathcal{S}^{(i-1)}$. Finally, we select top $N$ shots as candidates for further investigation, $\mathcal{C}^{(i)}=\{\mathcal{C}_{1}^{(i)},...\mathcal{C}_{N}^{(i)}\}$, where $\mathcal{C}_{n}^{(i)}$ refers to the $n$-th selected shot from the shot set $\mathcal{S}^{(i-1)}$. Additionally, since there is one single shot (i.e., the whole video) at the 1st round, we directly use this shot as the candidate at this round.
Shot Partition
Figure 3: Shot Partition. Given a candidate shot at step $i$, VideoChat-A1 first applies K-Means to obtain K cluster centers for finding key frames. Subsequently, subshots are partitioned based on the feature distance and its adjacent two key frames.
After obtaining the candidate shots $\mathcal{C}^{(i)}$, we next further divide each of them into subshots, to look deeper into it for reasoning. In this paper, we introduce a concise shot partition pipeline as shown in Fig. 3. For notation simplicity, we illustrate how to perform it on a candidate shot.
Key Frame Discovery. To discover the subshots in a candidate shot, we start by finding key frames in the subshots. First, we uniformly sample frames from the candidate shot. To reduce temporal redundancy while maintaining computation efficiency, we sample 1 fps in this paper. Second, we extract the features of the sampled frames by CLIP. Third, we perform K-Means (MacQueen 1967) on these features. As a result, each of $K$ clusters roughly reflects a subshot. The choice of $K$ is discussed in detail in the ablation study section. However, frames in each cluster may not be temporally adjacent, since K-Means Clustering does not take temporal order into account. To construct subshots in temporal order, we further find a key frame in each cluster to divide the shot. Specifically, for each cluster, we choose the frame whose feature is closest to the cluster center, as the key frame in this cluster. Finally, we organize $K$ key frames in the temporal order, $\mathcal{K}=\{\mathcal{K}(1),...,\mathcal{K}(K)\}$, for the candidate shot.
@@
Shot Reflection
Question Reasoning. After obtaining subshots from the candidate shots, we employ MLLMs to answer the user question based on these finer shot regions. Specifically, as shown in Fig. 4, we leverage MLLM as a dialogue agent to answer the question at round $i$,
@@
where $\mathcal{Q}$ and $\mathcal{O}$ refer to question and answer options. We sample $\mathcal{T}$ frames from each subshot $\mathcal{S}_{n}^{(i)}(k)$, where $\mathcal{T}$ is the number of sampled frames, and combine them with $\mathcal{V}^{(i-1)}$ to form $\mathcal{V}^{(i)}$. Finally, given these inputs, MLLM generates answer $\mathcal{A}^{(i)}$, and the reason why to choose this answer $\mathcal{R}^{(i)}$.
@@
We set the confidence level $\mathcal{Z}^{(i)}$ from 0 to 3. If the confidence is higher than 2, we believe that it is a reliable answer as the final output. Otherwise, we believe that it is an unreliable answer and conduct further investigation. Hence, we start shot selection, partition, reflection in the next round. Additionally, we set a maximum number of shot reasoning rounds. If VideoChat-A1 still does not get the confident answer at the max step, it will vote the major answer as the final one, according to the answers of all the rounds. Via exploit relevant shots in a coarse-to-fine manner, our VideoChat-A1 allows to leverage a chain of shots for deep thinking on long videos, and thus progressively and interactively discover the reliable shot context to answer user question.
@@
Closed-Source Model
GPT-4o(0513) (OpenAI 2024) 72.2 66.7 64.6 80.0 / 82.8 70.3 / 76.6 65.3 / 72.1 71.9 / 77.2
Gemini 1.5 Pro (Reid et al. 2024) 71.1 64.0 - 81.7 / 84.5 74.3 / 81.0 67.4 / 77.4 75.0 / 81.3
@@
LLaVA-OneVision-72B (Li et al. 2024a) 62.0 63.2 68.0 76.7/79.3 62.2/66.9 60.0/62.4 66.3/69.6
VideoLLaMA-2-72B (Cheng et al. 2024) 63.9 - 45.6 69.8/72.0 59.9/63.0 57.6/59.0 62.4/64.7
LLAVA-Video-72B (Zhang et al. 2024b) 65.6 64.9 - 81.4/82.8 68.9/75.6 61.5/72.5 70.6/76.9
Qwen2.5-VL-72B (Bai et al. 2025)  77.9 - - 80.1/82.2 71.3/76.8 62.2/74.3 71.2/77.8
InternVL-2.5-78B (Chen et al. 2024b) - 63.6 75.7 82.8/83.2 70.9/74.1 62.6/64.8 72.1/74.0
InternVL-3-78B (Zhu et al. 2025) - 65.7 79.5 - - - 72.7 /75.7
@@
ShareGPT4Video-8B (Chen et al. 2024a) - 39.7 46.4 48.3/53.6 36.3/39.3 35.0/37.9 39.9/43.6
VideoChat2-7B (Li et al. 2024c) 56.7 39.3 47.9 48.3/52.8 37.0/39.4 33.2/39.2 39.5/43.8
LongVU-7B (Shen et al. 2024) 67.6 - 65.4 - - -/59.5 -/60.6
LLaVA-Video-7B (Zhang et al. 2024b) 57.3 58.2 70.8 - - - 63.3/69.7
Qwen2.5-VL-7B (Bai et al. 2025) 65.0 56.0 - - - - 65.1/71.6
InternVideo-2.5-8B (Wang et al. 2025) 63.9 60.6 72.8 - - - 65.1/-
InternVL-2.5-8B (Chen et al. 2024b) - 60.0 68.9 - - - 64.2/66.9
InternVL-3-8B (Zhu et al. 2025) - 58.8 71.4 - - - 66.3/68.9
@@
VideoAgent (Wang et al. 2024a) 54.1 - - - - - -
VideoTree (GPT-4)  (Wang et al. 2024c) 61.1 - - - - - -
DrVideo (Ma et al. 2024) 61.0 - - - - - 51.7/71.7
T* (Ye et al. 2025) - - - 61.0/- 66.6/- 77.5/- 68.3/-
VideoMind (Liu et al. 2025b) - - 64.4 - - 49.2/- 58.2/-
VideoMultiAgents (Kugo et al. 2025) 68.0 - - - - - -
VideoRAG-72B (Luo et al. 2024) - 65.4 73.8 81.1/- 72.9/- 73.1/- 75.7/-
@@
Table 1: Comparison with closed-source, open-source, and agent based model on four Long Video Understanding Tasks. The Video-MME results are presented in the format “w/o subs / w/ subs".
@@
GPT-4o(OpenAI 2024) 384 134.4s 72.2 66.7 64.6 71.9/77.2
Gemini-1.5-Pro(Reid et al. 2024) 568 198.8s 71.1 64.0 - 75.0/81.3
Qwen2.5-VL-72B(Bai et al. 2025) 768 122.4s 77.9 - - 71.2/77.8
Qwen2.5-VL-7B(Bai et al. 2025) 512 24.3s 66.7 55.6 - 63.3/69.7
InternVL-2.5-8B(Chen et al. 2024b) 64 12.4s - 60.0 68.9 64.2/66.9
InternVideo-2.5-8B(Wang et al. 2025) 512 33.8s 63.9 60.6 72.8 65.1/-
@@
Table 2: Comparison on Average Frame Number and Inference Time on Datasets. Inference Time includes the average time for video frame extraction, clustering, and the time required for the MLLM to answer all queries.
Experiment
Experimental Setup